Fixed Assets and Depreciation

In recent years, Kevin Transportation purchased three used buses. Because of frequent
turnover in the accounting department, a different accountant selected the depreciation method
for each bus, and various methods were selected. Information concerning the buses is
summarized as follows.

Bus Acquired Cost Salvage Value Useful Iife in Years Depreciation Method

1 1/1/18 $96,000 6000 5 straight line

2 1/1/18 $110,000 10000 4 declining

3 1/1/19 $92,000 8000 5 units of activity

For the declining-balance method, the company uses the double-declining rate. For the
units-of-activity method, total miles are expected to be 120,000. Actual miles of use in the first 3
years were 2019, 24,000; 2020, 34,000; and 2021, 30,000. Instructions (a) Compute the amount
of accumulated depreciation on each bus at December 31, 2020.
Answer #1
Bus 1 Bus 2 Bus 3
Accumulated Depreciation at December 31, 2020 $ 54,000 $ 96,250 $ 40,600

Bus 1 :

Accumulated depreciation as on December 31, 2020 = [$ ( 96,000 - 6,000 ) / 5 ]* 3 = $ 54,000.

Bus 2 :

Double declining rate = (100 / 4) * 2 = 50 %

Depreciation expense for 2018 = $ 110,000 * 50 % = $ 55,000

Depreciation expense for 2019 = $ ( 110,000 - 55,000 ) * 50 % = $ 27,500

Depreciation expense for 2020 = $ ( 110,000 - 55,000 - 27,500) * 50 % = $ 13,750

Accumulated depreciation as on December 31, 2020 = $ ( 55,000 + 27,500 + 13,750 ) = $ 96,250.

Bus 3 :

Depreciation rate per mile = $ ( 92,000 - 8,000 ) / 120,000 miles = $ 0.70 per mile.

Accumulated depreciation as of December 31, 2020 = ( 24,000 + 34,000 ) * $ 0.70 = $ 40,600